Calculate the osmotic pressure of an aqueous solution containing 25.0 kg of sucrose (C₁₂H₂₂O₁₁) in 0.50 m³ of a solution at 300 K . (Molar mass of sucrose is 342 kɡ mol ⁻¹ .)
1
Expert's answer
2019-11-18T06:28:42-0500

Osmotic pressure is calculated by the formula:

p=cRT, where

c - molar concentration (mol/m3);

R - universal gas constant = 8,314 J/mol•K;

T - temperature, K.

To calculate the concentration, we use this formula:

c = m/(M•V) = 25/(342•0,5) = 0,146 mol/m3.

So, p = 0,146•8,314•300 = 364,65 J/m3 = 364,65 Pa